# How Beginning Programmers and Code LLMs (Mis)read Each Other

### v1

@@ Abstract.
Figure 1. Visualization of the multi-step process of querying an large language mode of code (Code LLM). The user starts with crafting their prompt in natural language (NL). They provide the prompt to the model, which produces code. The user then assesses the correctness of the generated code. If there are errors, they must identify how to resolve them and how to edit the prompt. This continues in an iterative fashion.
@@ 1. Introduction
We find that students struggle significantly with this task, even though we pose problems tailored to their skill level and test code correctness for them. In essence, beginning programmers and current Code LLMs tend to misread each other: the Code LLM fails to generate working code based on student descriptions and students have a hard time adapting their descriptions to the model. Our study has concerning implications for democratizing programming: if these students, who already have basic skills in code explanation and understanding, struggle with this simplified task, the full natural language-to-code task—where the user has to determine correctness themselves—must be very challenging indeed for true novices. This finding also has important implications for education. Code LLMs have sparked an intense debate over the future of computing education, including claims that traditional programming training is no longer necessary (Welsh, 2022; Manjoo, [n. d.]). By contrast, our findings highlight the continuing importance of teaching students technical communication and code understanding.
Our work differentiates itself from previous work in three key ways: scale, population, and experimental design. First, we study 120 students solving 48 different programming problems. To our knowledge, no previous work has studied user interactions with Code LLMs at this scale. Second, we focus on a near-novice population with fairly uniform levels of experience, allowing us to carefully tailor tasks to their skill level. Finally, we use an experimental paradigm that allows us to isolate the prompt writing and editing aspects of the task.
@@ 2. Related Work
Our work focuses on how programmers use LLMs to turn natural language into code. Programming with natural language is a decades old idea (Miller, 1981) and has led to several ideas about bringing programming closer to how users communicate (Myers et al., 2016). For instance, Hindle et al. (2012) imagined that future language models could be effective at turning natural language to code, a prediction that has been borne out with Code LLMs.
@@
Like us, several researchers have considered the impact of using Code LLM for the text-to-code task with non-experts, specifically in educational settings. Our work is larger in scale than prior work (120 students from 3 institutions and 48 problems in 8 categories), which allows us to perform statistical analyses that require large sample sizes to be reliable. Moreover, our experiment design allows us to investigate key research questions that prior work has not been able to ask, such as identifying the prompting strategies that beginners use, determining how they modify prompts that do not work, and studying several factors that affect their success.
@@
Lau and Guo (2023) interviewed 20 CS1/CS2 instructors in early 2023 about their perceptions of ChatGPT and LLM technologies. They report that instructors hold a diverse set of perspectives: some wanted to “ban it” and others felt urged to integrate these technologies into curricula to prepare students for future jobs may require using LLM technology. The students in our study echo many of the concerns and desires raised by instructors in Lau and Guo (2023).
@@
Copilot and related tools suggest inline code completions, but there are other ways to interact with AI-assisted programming tools. Vaithilingam et al. (2023) present new interfaces for Visual Studio that present code changes. Liu et al. (2023b) build a new interaction model, grounded abstraction matching (Liu et al., 2023b), which targets spreadsheets and data frames, constraining the generated code to support grounding. These ideas are exciting parallel directions for Code LLM interaction in addition to the natural language prompting approach we study here.
@@
Recent additional efforts include Finnie-Ansley et al. (2022), who report that Codex is remarkably good at generating code from natural language prompts from a CS1 class and several variations of the Rainfall Problem; Dakhel et al. (2022), who compare the quality of Codex-generated code to student-written code; and Babe et al. (2023), who use student-written prompts to benchmark Code LLMs. Finally, Code LLMs have applications that go beyond natural-language-to-code, and researchers are using them as building blocks for a variety of other tasks (Schäfer et al., 2023; Bareiß et al., 2022; Murali et al., 2023; Nam et al., 2023; Ross et al., 2023; Lemieux et al., 2023; Xia et al., 2023; Joshi et al., 2023; Phung et al., 2023; First et al., 2023; Chen et al., 2023; Ferdowsi et al., 2023). The aforementioned papers present new tools, benchmarks, and studies of LLM capabilities. But, they do not study users’ abilities to prompt models, which is the focus of our work.
@@ 3. Study Design
Our study limits user interactions in order to isolate prompt writing and editing strategies. One key feature of our paradigm is that we automatically test the generated code. In most observational studies, programmers determine on their own whether the generated code is correct. This is itself an interesting process. However, studying this aspect of Code LLM interaction comes at the cost of studying prompt editing: if a programmer mistakenly accepts incorrect code, they will move on to the next task without editing. Prather et al. (2023) report that many of their participants mistakenly accepted incorrect code. Beginning students are particularly likely to err in this way: they may struggle to understand generated code, and their lack of confidence in their own abilities may make them trust the automated system over their own judgment (an example of automation bias (SKITKA et al., 2000; Goddard et al., 2012; Gadala, 2017; De-Arteaga et al., 2020)).
@@
We therefore rely on a popular alternative for describing program behavior: input/output examples (Figure 3). Students also could reference the function name and parameter names. Our participants had taken CS1 classes where natural language descriptions are typically accompanied by input/output examples, making this a familiar way of communicating program behavior. Several CS1 courses, and some of the assignments used in our CS1 courses, go beyond this and require students to construct their own examples or even practice test-driven development (Flatt et al., 2001; Edwards, 2004). However, our study does not require students to write their own tests.
Avoiding natural language presentation is critical in order to study how beginning programmers describe problems in their own words. However, it comes with two risks. First, the input/output paradigm may increase task difficulty, since participants must identify the key pattern on their own. Although understanding natural language descriptions of coding tasks is not always easy for beginning programmers, it is likely easier than our input/output paradigm. Second, input/output examples run the risk of underspecification (Shaw et al., 1975; Gulwani et al., 2017) – there may be more than one program that performs the correct input-output mapping. To determine that the provided tests adequately described the problem, we confirmed that our provided test sets had 100% code coverage for a correct solution and performed mutation testing (Jia and Harman, 2010). We also calculated participants’ success using only the provided test cases: if the generated code passed the provided tests, it was deemed correct, ensuring that the problem presentation aligned directly with the feedback to the user.
@@
Our main goal was to select problems at an appropriate level for students who had completed only CS1. Since our research questions focus on student prompting processes, not learning outcomes, we chose problems at a similar level to what participants might be able to code independently. Asking students to solve new or more complex problem types increases the likelihood that the Code LLM will generate unfamiliar or difficult to understand code, making the prompt editing process more difficult. We therefore adapted problems specifically from CS1 course materials at each institution. We made small changes to facilitate input/output testing or adjust problem difficulty.
We selected 48 problems balanced across eight conceptual categories from CS1 (Figure 2), similar to Kazemitabaar et al. (2023), but with more categories and problems. Each individual problem was assigned to 20 students; we balance the experimental lists to control for ordering effects, so that each participant solved one problem in each category, and the average difficulty of each problem list was roughly the same. To facilitate difficulty and category coverage, previous CS1 instructors were asked to provide additional problems as needed. Problems such as exp (Figure 3), for instance, require students to only recognize that numbers in a list are being squared. Other problems ask students to remember complex data structures (e.g. lists, dictionaries), but not the specific Python syntax for them. We further discuss student understanding of the problems in Section  7.2.
@@ 4. Study Logistics
(b) We run expert tests automatically and highlight ones they fail. Students are then able to either edit their description by pressing ”Try Again” or move on to another problem.
@@ 4.1. Charlie Interface
Charlie does not permit participants to edit the generated code. If the code fails, they can retry the problem or move ahead. For retry attempts, we pre-fill the text box with their last prompt to make editing easier. Finally, after every final attempt at a problem, Charlie presents two forced-choice questions with thumbs-up / thumbs-down answers: Did Charlie generate correct code? and Would you have written this code yourself?.
Each student worked with Codex to solve 3 tutorial problems and 8 main problems. We used the Charlie character to provide distance from any AI system that students might already know. This suggested a representation that was not human and not robotic. Charlie also provides visual feedback: Charlie animates a “thinking” position while Codex generates a completion and appears in different forms when the code does or does not pass all tests.
@@ 4.2. Model Choice
When we began piloting in November 2022, the most capable Code LLM was the largest Codex model from OpenAI, code-davinci-002. Although code-davinci-002 was first released in 2021, on established Python programming benchmarks, it remains as good as gpt-3.5-turbo, which is the model presently used for GitHub Copilot’s inline completions (Zhao, 2023), the free version of ChatGPT, and several other commercial products. Specifically, gpt-3.5-turbo and code-davinci-002 score 48% and 46% respectively on the HumanEval Python programming benchmark (OpenAI, 2023; Cassano et al., 2023), the most commonly used Python benchmark for Code LLMs. Since we started our study, several other LLMs have also appeared, including non-proprietary LLMs that are better for reproducibility (§ 9). The best open models perform comparably to code-davinci-002; for instance, CodeLlama (34B) achieves 48% on HumanEval (Rozière et al., 2023a). This suggests that the model that we use is as capable at code completion as newer models used in practice.
There are larger models that are more capable, such as GPT-4, which achieves a HumanEval score of 67% (OpenAI, 2023). However, GPT-4 is significantly slower and higher latency than the alternatives, and low latency is essential for LLM code completion to be acceptable to users (Murali et al., 2023); if participants have to wait more than a few seconds for the generated code, their frustration might lead them to move on rather than re-attempting the problem.
For consistency, we used the same Codex model (code-davinci-002) throughout the study. It is important to note that Code LLMs perform best when their output is sampled; consequently, the model may produce different programs for the same prompt. We generated output using best practices for hyperparameter and sampler settings (Chen et al., 2021).
@@ 4.3. Participants
We recruited 40 participants from each institution (n=120). Eligible participants were at least 18 years old, had taken CS1 at their institution between Fall 2021 and Spring 2023, and had not completed any subsequent CS courses. We recruited participants from March-July 2023 until reaching our sample size of 120. The pilot and main study received IRB approval.
@@
Figure 4. A diagram of the Charlie system including the front-end interface, the back-end addition of prompts to the function declaration, sending declaration + prompt to Codex via the API, running the code completion from Codex on our pre-defined tests, and returning the results to the participant in the front-end interface.
A diagram of the Charlie system including interface, back-end processing of input and connection to Codex, reprocessing data, and feedback to the user
Figure 4. A diagram of the Charlie system including the front-end interface, the back-end addition of prompts to the function declaration, sending declaration + prompt to Codex via the API, running the code completion from Codex on our pre-defined tests, and returning the results to the participant in the front-end interface.
@@ 4.4. Study Execution
The main experiment consisted of 8 problems in two blocks, the first untimed, the second timed. In the second block, students were limited to 5 minutes per problem. Participants were randomly assigned experimental lists, balanced by difficulty, using a modified Latin Square design. Four authors independently assessed each problem’s difficulty; we averaged these scores and developed six roughly equal lists (Figure 2).
After the main study, students completed a two-part survey, a semi-structured interview, and an optional debriefing session (Figure 2 (1)). The semi-structured interview was interleaved between two survey blocks to mitigate question ordering and priming biases. The first series of questions in the post-survey were adapted from previous work (Körber, 2018; Bartneck et al., 2009; Wang et al., 2021; Druga and Ko, 2021; Gorson and O’Rourke, 2020; Hart and Staveland, 1988) asking about usability, trust in automation, AI perception, growth mindset, and cognitive load. The semi-structured interview asked 8 questions covering student editing processes, what they found hard or easy, how they envisioned their interactions with Charlie, and how they imagined Charlie worked. In the optional debriefing, we explained the experiment and how Code LLMs work. The second part of the survey focused on participants’ backgrounds and demographics.

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642706isbn: 979-8-4007-0330-0/24/05conference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642706isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing User studiesccs: Social and professional topics Computing educationccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learningccs: Software and its engineering
Figure 1. Visualization of the multi-step process of querying a large language mode of code (Code LLM). The user starts with crafting their prompt in natural language (NL). They provide the prompt to the model, which produces code. The user then assesses the correctness of the generated code. If there are errors, they must identify how to resolve them and how to edit the prompt. This continues in an iterative fashion.
@@ 1. Introduction
We find that students struggle significantly with this task, even though we pose problems tailored to their skill level and test code correctness for them. In essence, beginning programmers and current Code LLMs tend to misread each other: the Code LLM fails to generate working code based on student descriptions and students have a hard time adapting their descriptions to the model. Our study has concerning implications for democratizing programming: if these students, who already have basic skills in code explanation and understanding, struggle with this simplified task, the full natural language-to-code task—where the user has to determine correctness themselves—must be very challenging indeed for true novices. This finding also has important implications for education. Code LLMs have sparked an intense debate over the future of computing education, including claims that traditional programming training is no longer necessary (Manjoo, 2023; Welsh, 2022). By contrast, our findings highlight the continuing importance of teaching students technical communication and code understanding.
Our work differentiates itself from previous work in three key ways: scale, population, and experimental design. First, we study 120 students solving 48 different programming problems. To our knowledge, no previous work has studied user interactions with Code LLMs at this scale. Second, we focus on a near-novice population with fairly uniform levels of experience, allowing us to carefully tailor tasks to their skill level. Finally, we use an experimental paradigm that allows us to isolate the prompt writing and editing aspects of the task. [^1] Data collected as part of this work is publicly available at https://doi.org/10.17605/OSF.IO/V2C4T.
@@ 2. Related Work
Our work focuses on how programmers use LLMs to turn natural language into code. Programming with natural language is a decades old proposition (Miller, 1981) and has led to several ideas about bringing programming closer to how users communicate (Myers et al., 2016). For instance, Hindle et al. (2012) imagined that future language models could be effective at turning natural language to code, a prediction that has been borne out with Code LLMs.
@@
Like us, several researchers have considered the impact of using Code LLMs for the text-to-code task with non-experts, specifically in educational settings. Our work is larger in scale than prior work (120 students from 3 institutions and 48 problems in 8 categories), which allows us to perform statistical analyses that require large sample sizes to be reliable. Moreover, our experiment design allows us to investigate key research questions that prior work has not been able to ask, such as identifying the prompting strategies that beginners use, determining how they modify prompts that do not work, and studying several factors that affect their success.
@@
Lau and Guo (2023) interviewed 20 CS1/CS2 instructors in early 2023 about their perceptions of ChatGPT and LLM technologies. They report that instructors hold a diverse set of perspectives: some wanted to “ban it” and others felt urged to integrate these technologies into curricula to prepare students for future jobs that may require using LLM technology. The students in our study echo many of the concerns and desires raised by instructors in Lau and Guo (2023).
@@
Copilot and related tools suggest inline code completions, but there are other ways to interact with AI-assisted programming tools. Vaithilingam et al. (2023) present new interfaces for Visual Studio that present code changes. Liu et al. (2023b) build a new interaction model, grounded abstraction matching, which targets spreadsheets and data frames, constraining the generated code to support grounding. These ideas are exciting parallel directions for Code LLM interaction in addition to the natural language prompting approach we study here.
@@
Recent additional efforts include Finnie-Ansley et al. (2022), who report that Codex is remarkably good at generating code from natural language prompts from a CS1 class and several variations of the Rainfall Problem; Dakhel et al. (2022), who compare the quality of Codex-generated code to student-written code; and Babe et al. (2023), who use student-written prompts to benchmark Code LLMs. Finally, Code LLMs have applications that go beyond natural-language-to-code, and researchers are using them as building blocks for a variety of other tasks (Schäfer et al., 2023; Bareiß et al., 2022; Murali et al., 2023; Nam et al., 2024; Ross et al., 2023; Lemieux et al., 2023; Xia et al., 2024; Joshi et al., 2023; Phung et al., 2023; First et al., 2023; Chen et al., 2023; Ferdowsi et al., 2023). The aforementioned papers present new tools, benchmarks, and studies of LLM capabilities. But, they do not study users’ abilities to prompt models, which is the focus of our work.
@@ 3. Study Design
Our study limits user interactions in order to isolate prompt writing and editing strategies. One key feature of our paradigm is that we automatically test the generated code. In most observational studies, programmers determine on their own whether the generated code is correct. This is itself an interesting process. However, studying this aspect of Code LLM interaction comes at the cost of studying prompt editing: if a programmer mistakenly accepts incorrect code, they will move on to the next task without editing. Prather et al. (2023) report that many of their participants mistakenly accepted incorrect code. Beginning students are particularly likely to err in this way: they may struggle to understand generated code, and their lack of confidence in their own abilities may make them trust the automated system over their own judgment (an example of automation bias (Skita et al., 2000; Goddard et al., 2012; Gadala, 2017; De-Arteaga et al., 2020)).
@@
We therefore rely on a popular alternative for describing program behavior: input/output examples (Figure 3). Students also could reference the function name and parameter names. Our participants had taken CS1 classes where natural language descriptions are frequently accompanied by input/output examples (see Appendix A.1.2), making this a familiar way of communicating program behavior. Several CS1 courses, and some of the assignments used in our CS1 courses, go beyond this and require students to construct their own examples or even practice test-driven development (Flatt et al., 2001; Edwards, 2004). However, our study does not require students to write their own tests.
Avoiding natural language presentation is critical in order to study how beginning programmers describe problems in their own words. However, it comes with two risks. First, the input/output paradigm may increase task difficulty, since participants must identify the key pattern on their own. Although understanding natural language descriptions of coding tasks is not always easy for beginning programmers, it is likely easier than our input/output paradigm. Second, input/output examples run the risk of underspecification (Shaw et al., 1975; Gulwani et al., 2017) – there may be more than one program that performs the correct input-output mapping. To determine that the provided tests adequately described the problem, we confirmed that our provided test sets had 100% code coverage for a correct solution and performed mutation testing (Jia and Harman, 2011). We also calculated participants’ success using only the provided test cases: if the generated code passed the provided tests, it was deemed correct, ensuring that the problem presentation aligned directly with the feedback to the user.
@@
Our main goal was to select problems at an appropriate level for students who had completed only CS1. Since our research questions focus on student prompting processes, not learning outcomes, we chose problems at a similar level to what participants might be able to code independently. Asking students to solve new or more complex problem types increases the likelihood that the Code LLM will generate unfamiliar or difficult to understand code, making the prompt editing process more difficult. We therefore adapted Python problems specifically from CS1 course materials at each institution. We made small changes to facilitate input/output testing or adjust problem difficulty. Appendix A.1 contains two examples of how source problems were adapted.
We selected 48 problems balanced across eight conceptual categories from CS1 (Figure 2), similar to Kazemitabaar et al. (2023), but with more categories and problems. Each individual problem was assigned to 20 students; we balanced the experimental lists to control for ordering effects, so that each participant solved one problem in each category, and the average difficulty of each problem list was roughly the same. To facilitate difficulty and category coverage, previous CS1 instructors were asked to provide additional problems as needed. Problems such as exp (Figure 3), for instance, require students to only recognize that numbers in a list are being squared. Other problems ask students to remember complex data structures (e.g. lists, dictionaries), but not the specific Python syntax for them. We further discuss student understanding of the problems in Section  7.2 and Appendix B.2.
@@ 4. Study Logistics
(b) We run expert tests automatically and highlight ones that fail. Students are then able to either edit their description by pressing ”Try Again” or move on to another problem.
@@ 4.1. Charlie Interface
Charlie does not permit participants to edit the generated code, since we are focused on natural-language-to-code interactions. If the code fails, they can retry the problem or move to the next problem. For retry attempts, we pre-fill the text box with their last prompt to make editing easier. Finally, after every final attempt at a problem, Charlie presents two forced-choice questions with thumbs-up / thumbs-down answers: Did Charlie generate correct code? and Would you have written this code yourself?. We included these questions to gather information about student perceptions of code style, since the model may produce working code, but in a style that is unfamiliar to students.
Each student worked with Codex to solve 3 tutorial problems and 8 main problems. We used the Charlie character to provide distance from any AI system that students might already know. This suggested a representation that was not human and not robotic. Charlie also provides visual feedback: Charlie animates a “thinking” position while Codex generates a completion and appears in different forms when the code does or does not pass all tests. We made this design choice to mitigate frustration with waiting for the model to generate code, a source of annoyance in prior studies of Code LLM interactions (Murali et al., 2023).
@@ 4.2. Model Choice
When we began piloting in November 2022, the most capable Code LLM was the largest Codex model from OpenAI, code-davinci-002. Although code-davinci-002 was first released in 2021, on established Python programming benchmarks, it remains as good as gpt-3.5-turbo, which is the model presently used for GitHub Copilot’s inline completions (Zhao, 2023), the free version of ChatGPT, and several other commercial products. Specifically, gpt-3.5-turbo and code-davinci-002 score 48% and 46% respectively on the HumanEval Python programming benchmark (OpenAI et al., 2023; Cassano et al., 2023), the most commonly used Python benchmark for Code LLMs. Since we started our study, several other LLMs have also appeared, including non-proprietary LLMs that are better for reproducibility (§ 9.5). The best open models perform comparably to code-davinci-002; for instance, CodeLlama (34B) achieves 48% on HumanEval (Rozière et al., 2024). This suggests that the model that we use is as capable at code completion as newer models used in practice.
There are larger models that are more capable, such as GPT-4, which achieves a HumanEval score of 67% (OpenAI et al., 2023). However, GPT-4 is significantly slower and higher latency than the alternatives, and low latency is essential for LLM code completion to be acceptable to users (Murali et al., 2023); if participants have to wait more than a few seconds for the generated code, their frustration might lead them to move on rather than re-attempting the problem.
For consistency, we used the same Codex model throughout the study (code-davinci-002). It is important to note that Code LLMs perform best when their output is sampled; consequently, the model may produce different programs for the same prompt. We generated output using best practices for hyperparameter and sampler settings (Chen et al., 2021).
@@ 4.3. Participants
We recruited 40 participants from each institution (n = 120). Eligible participants were at least 18 years old, had taken CS1 at their institution between Fall 2021 and Spring 2023, and had not completed any subsequent CS courses. We recruited participants from March to July 2023 until reaching our sample size of 120. The pilot and main study received IRB approval.
@@
Figure 4. An overview of the experimental platform. For each problem, the frontend provides the participant with the signature and tests and asks them to write a description (prompt). This is then relayed to the backend, where the signature and prompt are sent to Codex via the API. The code completion from Codex is then run on our pre-defined tests. Finally, the results of running the tests and the code completion are presented to the participant in the frontend interface.
[…]
A diagram of the Charlie system including interface, backend processing of input and connection to Codex, reprocessing data, and feedback to the user
@@ 4.4. Study Execution
The main experiment consisted of 8 problems in two blocks, the first untimed, the second timed. In the second block, students were limited to 5 minutes per problem. We included both timed and untimed blocks in order to balance the need to bound study duration with the desire to observe complete prompt editing cycles.
Participants were randomly assigned experimental lists, balanced by difficulty, using a modified Latin Square design. Four authors independently assessed the difficulty of writing prompts for each problem; we averaged these scores and developed six roughly equal lists (Figure 2).
Post-task Interview and Survey
After the main study, students completed a two-part survey, a semi-structured interview, and an optional debriefing session (Figure 2 (1)). The semi-structured interview was interleaved between two survey blocks to mitigate question ordering and priming biases.
The first part of the survey was designed to study student perceptions of Charlie and of AI more broadly. We adapted validated scales from previous work to understand student perceptions of the usability, trustworthiness, and friendliness of Charlie (Bartneck et al., 2009; Körber, 2018; Wang et al., 2021; Druga and Ko, 2021) and the mental workload of the task (Hart and Staveland, 1988). [^2] In some cases, we removed questions that were not relevant to our study to keep the survey length manageable for participants. Details available via our Supplemental Materials at https://doi.org/10.17605/OSF.IO/V2C4T. We were also interested in whether students’ ability to come up with effective prompting strategies might correlate with fixed versus growth mindsets about computing; we drew on Gorson and O’Rourke (2020) to measure this.
The semi-structured interview asked 8 questions covering student editing processes, what they found hard or easy, how they envisioned their interactions with Charlie, and how they imagined Charlie worked. The specific questions were directly inspired by our overarching research questions. Researchers followed a standing script to ask each question - there are a total of 5 missing question responses across the possible 960 interview datapoints, likely due to researcher error or time considerations. In the optional debriefing, we explained the experiment and how Code LLMs work.
The second part of the survey focused on participants’ backgrounds and demographics. These were the last questions of the study to mitigate possible stereotype threat (National Center for Women & Information Technology, 2023). For questions related to identity (e.g., gender, race, spoken language at home), we followed best practices and solicited responses via open text boxes (Spiel et al., 2019). We also asked questions about students’ CS1 performance, experience with programming outside of CS1, high school & educational background, math background, major, and class year.
Pilot Study